the mass of 4.0 x 10^23 formula units of potassium chloride (KCI)


1
Expert's answer
2021-10-12T07:48:26-0400

Solution:

One mole of any substance contains 6.022×1023 formula units (Na).

Hence,

Moles of KCl = (4.0×1023 formula units) × (1 mol / 6.022×1023 formula units) = 0.664 mol KCl


The molar mass of KCl is 74.55 g/mol

Hence,

Mass of KCl = (0.664 mol KCl) × (74.55 g KCl / 1 mol KCl) = 49.5 g KCl

Mass of KCl = 49.5 g


Answer: 49.5 grams